# Enhancing Delta Compression in LLMs via SVD-based Quantization Error Minimization

### v2

@@ Abstract
Fine-tuning is a crucial process for adapting large language models (LLMs) to diverse applications. In certain scenarios, like multi-tenant serving, a large number of LLMs finetuned from the same base model are deployed to meet complex requirements for users. Recent works explore delta-compression approaches to quantize and compress the delta weights between the customized LLM and the corresponding base model. However, they exhibit inadequate performance at high compression ratios due to their empirical nature. In this work, we introduce DeltaMix, an adaptive mixed-precision delta-compression framework designed to minimize quantization error in the singular value decomposition (SVD) space without imposing additional assumptions. DeltaMix provides a theoretical justification for the necessity of mixed-precision compression and presents a practical quantization solution that involves solving a 0/1 linear integer programming problem alongside a reconstruction target correction method. Experimental results across multiple models and benchmarks illustrate that DeltaMix consistently outperforms all baseline methods. Notably, on tasks such as AIME2024 and GQA, DeltaMix exceeds the performance of the best baseline, Delta-CoMe, by 22.3% and 6.1% for 7B parameter models, respectively.
@@ 1 Introduction
In this work, we propose DeltaMix, a high-performance mixed-precision delta-compression framework grounded in a solid theoretical foundation. DeltaMix implements delta-compression within the SVD space, formulating the quantization objective as the minimization of layer-wise quantization error. By pursuing this objective, DeltaMix establishes a mathematically sound mixed-precision compression strategy that accommodates flexible, user-defined compression ratios. This strategy derives the mixed-precision quantization scheme through the solution of a 0/1 linear integer programming problem and ensures optimization consistency throughout the quantization process via a reconstruction target correction method. Unlike Ping et al. (2024), which empirically posits that singular vectors corresponding to larger singular values are more significant and, therefore, necessitate higher-bit representations, DeltaMix prioritizes the minimization of quantization error. It formulates all subsequent strategies based exclusively on this principle, eschewing reliance on singular values for assessing importance. This distinction is vital, as prior research has demonstrated that the significance attributed to singular values may not correlate with the performance of LLMs (Hsu et al., 2022; Wang et al., 2025).
We conduct extensive experiments on reasoning, math, code, and multimodal tasks across eight aligned LLMs to demonstrate the effectiveness of DeltaMix. The results show that DeltaMix achieves state-of-the-art performance among delta-compression methods, particularly in challenging scenarios where the norm of $\Delta\mathbf{W}$ is large. Notably, on the reasoning task AIME2024, DeltaMix surpasses the leading baseline, Delta-CoMe, by 22.3% on the 7B model and 26.9% on the 14B model. Furthermore, DeltaMix can achieve more than 6× GPU memory and disk storage savings, enabling the deployment of multiple models within constrained resource environments.
Figure 1: An overview of DeltaMix. The quantization error of the $i$-th row of $\mathbf{V}$ comprise two components: a “scaling” term (➃) and a “difference” term (➂). The “scaling” term is fixed, while the “difference” term is related to the mixed-precision quantization scheme (➀). DeltaMix identifies the optimal quantization scheme within the constraints of the bit budget (➀) to effectively balance these two components, thereby minimizing the total quantization error of $\mathbf{V}$ (➄). Note that the "difference" term for various bit-widths (➁) is pre-computed using a calibration dataset and remains fixed during the optimization process.
@@ 2 Related Work
Delta-compression (Isik et al., 2023; Ryu et al., 2023; Liu et al., 2024; Ping et al., 2024) aims to diminish the storage and inference costs associated with serving multiple models by compressing delta parameters, which are the differences between the parameters of a fine-tuned LLM and its corresponding base LLM. GPT-Zip (Isik et al., 2023) extends GPTQ to compress the delta parameters into 2-bit, and then sparsify 95% of the quantized delta weights to further reduce storage costs. DeltaZip (Yao et al., 2024) extends the idea of structured pruning and delta-compression to develop a multi-tenant serving system. However, both methods are still limited to compression ratios of 2-bit and higher. Liu et al. (2024) introduces BitDelta, which compresses delta weight into 1-bit, using a trainable high-precision scaling factor for each delta weight matrix. From this point onward, the compression of delta parameters has entered the 1-bit era. In addition to these low-bit methods, Ryu et al. (2023) identifies the low-rank property of delta weights and achieves delta-compression through low-rank approximation. Recently, Delta-CoMe (Ping et al., 2024) leverages the benefits of both low-rank and low-bit compression methods, proposing a mixed-precision delta-compression method that uses varying bit-widths to represent different singular vectors of the delta weights. However, the rationale behind their mixed-precision quantization is predicated on a questionable hypothesis (Hsu et al., 2022; Wang et al., 2025): that singular vectors associated with larger singular values are inherently more important. This premise lacks a solid theoretical foundation, leading to a mixed-precision strategy that is primarily empirical and, consequently, suboptimal. In this work, we introduce DeltaMix, which provides a mathematical proof of the necessity for mixed-precision in SVD-based delta-compression methods, and derives a quantization approach that is firmly grounded in mathematical theory.
@@ 3 Method
Algorithm 1 Algorithm for Quantization in DeltaMix
In this section, we introduce DeltaMix, an adaptive mixed-precision delta-compression strategy for LLMs with mathematical support. In Section 3.1, we begin with the minimization of quantization error in the SVD space and derive the detailed quantization process. We provide a mathematical proof demonstrating the necessity of mixed-precision in this context. In Section 3.2, we introduce our mixed-precision schedule in detail, which is built on the solution of a 0/1 integer linear programming problem. Algorithm 1 shows the details of DeltaMix.
@@ 3.1 Quantization Error Derivation
At a high level, DeltaMix follows the structure of the classical post-training quantization method GPTQ, by performing quantization to minimize the reconstruction error. Given a delta weight matrix $\mathbf{W}$ and the corresponding input $X$, the quantization objective of the GPTQ is to find a quantized matrix $\mathbf{\hat{W}}$ which minimizes the squared error:
$\arg\min_{\mathbf{\hat{W}}}\left\|\mathbf{W}X-\mathbf{\hat{W}}X\right\|_{F}^{2}=\sum_{i}\left\|W_{i}X-\hat{W}_{i}X\right\|_{F}^{2}\approx\sum_{i}e_{i}$ (1)
@@
Instead of directly quantizing $\mathbf{W}$, DeltaMix performs quantization in the SVD space, by finding a quantized matrix $\mathbf{\hat{U}}$ and $\mathbf{\hat{V}}$ which minimizes the squared error:
@@
where $\mathbf{W}=\mathbf{U\Sigma V}$. Below, we introduce the detailed quantization process of DeltaMix, which first quantizes $\mathbf{V}$, and then moves to $\mathbf{U}$.
@@ 3.1.1 Quantize $\mathbf{V}$
Here $\Delta V_{i}=V_{i}-\hat{V}_{i}$ is the quantization difference of the $i^{\mathrm{th}}$ row, and $\mathbf{H}_{i}^{\mathbf{V}}=2\Sigma_{ii}^{2}\cdot XX^{T}$ is the Hessian matrix of the $i^{\mathrm{th}}$ row of $\mathbf{V}$ (with derivation details in Appendix C.1). As $\Sigma_{ii}^{2}$ is a scalar, we can reformulate the Eq. (4) as follows:
@@ 3.1.2 Quantize $\mathbf{U}$
Here $\Delta U_{i}=U_{i}-\hat{U}_{i}$, and the Hessian matrix of the $i^{\mathrm{th}}$ row of $\mathbf{U}$ is given by $\mathbf{H}^{\mathbf{U}}_{i}=2\mathbf{\Sigma\hat{V}}XX^{T}\mathbf{\hat{V}^{T}\Sigma^{T}}$(with derivation details in Appendix C.2). Upon comparing Eq. (5) and Eq. (6), we observe that $e_{i}^{\mathbf{U}}$ does not incorporate the scaling term present in Eq. (6). Consequently, when different rows are quantized using the same bit-width, there is no significant variation in error. This uniformity arises from the fact that the Hessian matrices for different rows of $\mathbf{U}$ are identical. Thus, unlike $\mathbf{V}$, there is no necessity to employ mixed precision when quantizing different rows of $\mathbf{U}$.
Therefore, DeltaMix determines the mixed-precision quantization schedule based on $\mathbf{V}$, and then applies the same schedule to $\mathbf{U}$ for simplicity. Specifically, DeltaMix quantizes $\mathbf{U}$ using a column-wise mixed-precision schedule, where the $i^{\mathrm{th}}$ column of $\mathbf{U}$ adopts the same bit-width as the $i^{\mathrm{th}}$ row of $\mathbf{V}$ as they correspond to the same singular value. Notably, DeltaMix exhibits insensitivity to column-wise precision schedules, since GPTQ compensates for quantization-induced errors in the column direction by adjusting the unquantized weights during the quantization process. This compensation, however, does not occur between different rows, as different rows are independently quantized in GPTQ. This further underscores the importance of discussing row-wise mixed precision strategies aimed at minimizing the quantization error of $\mathbf{V}$. In Appendix E.1, we further demonstrate experimentally that applying the same mixed-precision quantization strategy to both $\mathbf{V}$ and $\mathbf{U}$ yields satisfactory performance.
@@
In Eq. (6), we quantize $\mathbf{U}$ to reconstruct the target $\mathbf{U\Sigma}\hat{\mathbf{V}}X$, which deviates from the initial target $\mathbf{U\Sigma V}X$. This deviation can negatively impact the performance of the quantized model. A straightforward approach to address this issue is to directly replace the reconstruction target with $\mathbf{U\Sigma V}X$; however, this would inhibit the application of GPTQ for quantization since it would no longer rely exclusively on the Hessian matrix $\mathbf{H}^{\mathbf{U}}$. Therefore, we propose a method termed “Reconstruction Target Correction” (RTC) to reduce the bias by transforming $\mathbf{U\Sigma\hat{V}}X$ in Eq. (6) to $\mathbf{\tilde{U}\Sigma\hat{V}}X$, where $\mathbf{\tilde{U}}$ is derived from the following equation:
@@
See Appendix C.3 for detailed derivations. In summary, prior to quantizing $\mathbf{U}$, we first update $\mathbf{U}$ to $\mathbf{\tilde{U}}$ using Eq. (7). Subsequently, we perform quantization by minimizing $\|\mathbf{\tilde{U}\Sigma\hat{V}}X-\mathbf{\hat{U}\Sigma\hat{V}}X\|_{F}^{2}$. This approach aims to ensure that the reconstruction target closely approximates the original, without compromising the application of GPTQ for quantization.
@@ 3.2 Optimization Problem Modeling
$\min_{\mathcal{S}}\sum_{i}\mathbb{E}^{\mathbf{V}}_{i}\mathcal{S}_{i}^{T}$ (Total quantization error) (8)
s.t. $\sum_{i}\mathcal{S}_{i}B\leq G_{b}(h_{\mathrm{in}}\cdot h_{\mathrm{out}})$ (Bit budget constraint)
$\mathrm{sum}(S_{i})=1$ (One-hot vector constraint)
$S_{i}-f\leq 0$ (Bit-width selection constraint)
$\mathrm{sum}(f)\leq f_{\text{max}}$ (Bit-width number constraint)
@@
The $0/1$ integer linear programming optimization problem is then solved with the CVXPY (Diamond & Boyd, 2016) library and the SCIP (Maher et al., 2016) solver. We report the optimization solving time in Appendix E.3, which costs 29.4 minutes for Qwen2.5-Math-7B-Instruct. This overhead is acceptable, as the model requires quantization only once. By solving Eq. (8), we obtain an optimal mixed-precision quantization scheme that minimizes the error while satisfying predefined bit budget constraints. This allows us to derive task-specific mixed-precision quantization strategies which balance the “scaling” and “difference” terms, leading to improved performance across various tasks.
4 Experiments
4.1 Experiment Setup
Evaluation Tasks
We evaluate our methods on four distinct tasks: reasoning, math, code generation, and multi-modal. These tasks encompass a vast array of current directions based on fine-tuning with open-source LLMs. Reasoning: We use the Math500 and AIME2024 datasets as the test set. Math: We use the GSM8K  (Cobbe et al., 2021) and Math500 (Lightman et al., 2023) datasets as the test set. Code Generation: We use HumanEval  (Chen et al., 2021) and MBPP  (Austin et al., 2021) as the test set. Multi-Modal: We utilize the GQA  (Hudson & Manning, 2019) and the image part of ScienceQA (Lu et al., 2022) datasets. Please refer to Appendix D.1 for more details.
Models
To ensure a comprehensive comparison, we evaluate both 7B and 13-14B models across the four tasks with various backbones. See Table 7 in Appendix D.1 for more details about the backbones and aligned models used. During inference, we employ a greedy search strategy.
Baselines
We compare DeltaMix with three baselines: SVD-based low-rank compression (Ryu et al., 2023), BitDelta  (Liu et al., 2024), and Delta-CoMe  (Ping et al., 2024). All methods are evaluated using NVIDIA L20 GPUs.
4.2 Main Results
Table 1: Comparison of DeltaMix and baselines on various tasks across 7B-sized models. We report the results in the format “mean(std)” with three runs for Delta-CoMe and DeltaMix.
@@ 4.2 Main Results
DeltaMix 1/16 82.7(0.83) 36.7(3.35) 77.7(1.03) 94.6(0.51) 85.6(0.35) 83.1(0.25) 52.4(2.30) 79.4(0.83) 74.0
Table 2: Comparison of DeltaMix and baselines on various tasks across 13-14B-sized models. We report the results in the format “mean(std)” with three runs for Delta-CoMe and DeltaMix.
@@
DeltaMix 1/16 80.2(2.09) 31.1(3.81) 21.7(0.64) 71.2(0.26) 91.5(0.60) 86.9(0.12) 62.7(0.04) 72.1(0.18) 64.7
Tables 1 and 2 present the results of DeltaMix on both the 7B and 13-14B models across four tasks, in comparison to the baselines. Notably, DeltaMix demonstrates superior overall performance on both the 7B and 13-14B models, surpassing the best baseline, Delta-CoMe, by an average of 2.9% and 2.2%, respectively.
When analyzing the various tasks, we observe that DeltaMix exhibits more pronounced improvements in challenging scenarios characterized by a significant performance gap between the baseline methods and the aligned model. This is particularly evident in reasoning-intensive benchmarks, such as AIME2024, as well as in multimodal tasks utilizing 7B backbones. For instance, DeltaMix surpasses the previous state-of-the-art model, Delta-CoMe, by 22.3% on the 7B model and by 26.9% on the 14B model. Further analysis reveals that these models display larger norms for $\Delta\mathbf{W}$. Specifically, the median norm of DeepSeek-R1-Distill-Qwen-7B and Qwen2.5-VL-Instruct is 6.5 and 10.3 times that of Qwen-Coder-Instruct-7B, with corresponding values of 26.13 and 41.45 compared to 4.02, respectively. In this context, baseline methods struggle to achieve optimal solutions due to their empirical nature. In contrast, DeltaMix directly optimizes quantization error from a mathematical perspective, enabling it to fully leverage its strengths in demanding tasks. However, on tasks where baselines already achieve near-lossless accuracy, such as MBPP and HumanEval on the 7B backbone, DeltaMix performs comparably to the best baseline. In these scenarios, the norm of $\Delta\mathbf{W}$ is relatively small and can be easily compressed, leading to a ceiling effect: $\Delta\mathbf{W}$ can be quantized almost losslessly by existing baselines, leaving little room for further improvement.
We also compare the quantization time cost of DeltaMix and Delta-CoMe. Please refer to Appendix E.3 for more details. The results show that DeltaMix (resp. Delta-CoMe) requires only 1.2 (resp. 0.4)hours for 7B models and 2.4 (resp. 0.8) hours for 14B models on a single GPU. Although DeltaMix is slower than Delta-CoMe, the time cost remains acceptable since the quantification process needs to be performed only once.
4.3 Analysis of $f_{\mathrm{max}}$

### v3

# Principled SVD-based Delta Compression via Quantization Error Minimization

@@ Abstract
Supervised Fine-Tuning (SFT) empowers Large Language Models (LLMs) with exceptional performance on specialized tasks, but it yields dense, high-dimensional delta parameters that pose severe storage and distribution challenges. Singular Value Decomposition (SVD)-based compression offers a compact representation for such delta parameters, but existing methods adopt heuristic quantization without clarifying underlying mechanisms, leading to poor generalizability. In this work, we propose PrinMix, a rigorous SVD-based framework that models quantization as an optimization problem, grounding the design in mathematical mechanisms. We first theoretically derive quantization error and identify a key singular-value-dominated scaling mechanism, which mathematically proves the necessity of mix-precision quantization. We then model the quantization scheme as a 0/1 Integer Linear Programming (ILP) problem, which yields optimal bit-budget-constrained solutions without empirical assumptions. Furthermore, PrinMix integrates a Reconstruction Target Correction (RTC) method to compensate for errors from the $\mathbf{V}$-then-$\mathbf{U}$ sequential quantization process. Extensive experiments confirm PrinMix performs well: for 7B LLMs, PrinMix outperforms SOTA Delta-CoMe on challenging benchmarks by 22.3% on AIME2024 and 6.1% on GQA.
Delta-Compression, Post-Training Quantization Singular Value Decomposition,
@@ 1 Introduction
Figure 1: An overview of PrinMix compared to single-precision quantization. Given the quantization scheme (➀), we compute the “difference” term (➂) by looking up the corresponding values in subfigure ➁. The quantization error of the $i$-th row of $\mathbf{V}$ comprise two components: a “scaling” term (➃) and a “difference” term (➂). PrinMix identifies the optimal quantization scheme within the constraints of the bit budget (➀) to effectively balance these two components, thereby minimizing the total quantization error of $\mathbf{V}$ (➄). Note that the "difference" term for various bit-widths (➁) is pre-computed using a calibration dataset and remains fixed during the optimization process.
In this work, we propose PrinMix, a high-performance principled mixed-precision delta-compression framework grounded in a solid theoretical foundation. PrinMix implements delta-compression within the SVD space, formulating the quantization objective as the minimization of layer-wise quantization error. By pursuing this objective, PrinMix establishes a mathematically sound mixed-precision compression strategy that accommodates flexible, user-defined compression ratios. This strategy derives the mixed-precision scheme through the solution of a 0/1 linear integer programming problem and ensures optimization consistency throughout the quantization process via a reconstruction target correction method. Unlike Ping et al. (2024), which empirically posits that singular vectors corresponding to larger singular values are more significant and, therefore, necessitate higher-bit representations, PrinMix prioritizes the minimization of quantization error. It formulates all subsequent strategies based exclusively on this principle, eschewing reliance on singular values for assessing importance. This distinction is vital, as prior research has demonstrated that the significance attributed to singular values may not correlate with the performance of LLMs (Hsu et al., 2022; Wang et al., 2025).
We conduct extensive experiments on reasoning, math, code, and multimodal tasks across eight aligned LLMs to demonstrate the effectiveness of PrinMix. The results show that PrinMix achieves state-of-the-art performance among delta-compression methods, particularly in challenging scenarios where the norm of $\Delta\mathbf{W}$ is large. Notably, on the reasoning task AIME2024, PrinMix surpasses the leading baseline, Delta-CoMe, by 22.3% on the 7B model and 26.9% on the 14B model. Furthermore, PrinMix can achieve more than 6× GPU memory and disk storage savings, enabling the deployment of multiple models within constrained resource environments.
@@ 2 Related Work
Delta-compression (Isik et al., 2023; Ryu et al., 2023; Liu et al., 2024; Ping et al., 2024) aims to diminish the storage and inference costs associated with serving multiple models by compressing delta parameters, which are the differences between the parameters of a fine-tuned LLM and its corresponding base LLM. GPT-Zip (Isik et al., 2023) extends GPTQ to compress the delta parameters into 2-bit, and then sparsify 95% of the quantized delta weights to further reduce storage costs. DeltaZip (Yao et al., 2024) extends the idea of structured pruning and delta-compression to develop a multi-tenant serving system. However, both methods are still limited to compression ratios of 2-bit and higher. Liu et al. (2024) introduces BitDelta, which compresses delta weight into 1-bit, using a trainable high-precision scaling factor for each delta weight matrix. From this point onward, the compression of delta parameters has entered the 1-bit era. In addition to these low-bit methods, Ryu et al. (2023) identifies the low-rank property of delta weights and achieves delta-compression through low-rank approximation. Recently, Delta-CoMe (Ping et al., 2024) leverages the benefits of both low-rank and low-bit compression methods, proposing a mixed-precision delta-compression method that uses varying bit-widths to represent different singular vectors of the delta weights. However, the rationale behind their mixed-precision quantization is predicated on a questionable hypothesis (Hsu et al., 2022; Wang et al., 2025): that singular vectors associated with larger singular values are inherently more important. This premise lacks a solid theoretical foundation, leading to a mixed-precision strategy that is primarily empirical and, consequently, suboptimal. In this work, we introduce PrinMix, which provides a mathematical proof of the necessity for mixed-precision in SVD-based delta-compression methods, and derives a quantization approach that is firmly grounded in mathematical theory.
@@ 3 Method
Algorithm 1 Algorithm for Quantization in PrinMix
In this section, we introduce PrinMix, an adaptive mixed-precision delta-compression strategy for LLMs with mathematical support. In Section 3.1, we begin with the minimization of quantization error in the SVD space and derive the detailed $\mathbf{V}$-then-$\mathbf{U}$ sequential quantization process. We provide a mathematical proof demonstrating the necessity of mixed-precision in this context. In Section 3.2, we present the detailed design of our mixed-precision scheme, which is formulated as a 0-1 Integer Linear Programming (ILP) problem. Algorithm 1 shows the details of PrinMix.
@@ 3.1 Quantization Error Derivation
At a high level, PrinMix follows the structure of the classical post-training quantization method GPTQ, by performing quantization to minimize the reconstruction error. Given a delta weight matrix $\mathbf{W}$ and the corresponding input $X$, the quantization objective of the GPTQ is to find a quantized matrix $\mathbf{\hat{W}}$ which minimizes the squared error:
$\arg\min_{\mathbf{\hat{W}}}\left\|\mathbf{W}X-\mathbf{\hat{W}}X\right\|_{F}^{2}$ $=\sum_{i}\left\|W_{i}X-\hat{W}_{i}X\right\|_{F}^{2}$ (1)
$\approx\sum_{i}e_{i}$
@@
Instead of directly quantizing $\mathbf{W}$, PrinMix performs quantization in the SVD space, by finding a quantized matrix $\mathbf{\hat{U}}$ and $\mathbf{\hat{V}}$ which minimizes the squared error:
@@
where $\mathbf{W}=\mathbf{U\Sigma V}$. Below, we introduce the detailed $\mathbf{V}$-then-$\mathbf{U}$ sequential quantization process of PrinMix, which first quantizes $\mathbf{V}$, and then moves to $\mathbf{U}$.
@@ 3.1.1 Quantize $\mathbf{V}$
Here $\Delta V_{i}=V_{i}-\hat{V}_{i}$ is the quantization difference of the $i^{\mathrm{th}}$ row, and $\mathbf{H}_{i}^{\mathbf{V}}=2\Sigma_{ii}^{2}\cdot XX^{T}$ is the Hessian matrix of the $i^{\mathrm{th}}$ row of $\mathbf{V}$ (with derivation details in Appendix A.1). As $\Sigma_{ii}^{2}$ is a scalar, we can reformulate the Eq. (4) as follows:
@@
Task 7B Models 13-14B Models
Backbone Aligned Backbone Aligned
Math Qwen2.5-Math Qwen2.5-Math-Instruct LLaMA2 MetaMath
Multi-Modal Qwen2.5 Qwen2.5-VL-Instruct LLaMA2 LLAVA-V1.5
@@ 3.1.2 Quantize $\mathbf{U}$
Here $\Delta U_{i}=U_{i}-\hat{U}_{i}$, and the Hessian matrix of the $i^{\mathrm{th}}$ row of $\mathbf{U}$ is given by $\mathbf{H}^{\mathbf{U}}_{i}=2\mathbf{\Sigma\hat{V}}XX^{T}\mathbf{\hat{V}^{T}\Sigma^{T}}$(with derivation details in Appendix A.2). Upon comparing Eq. (5) and Eq. (6), we observe that $e_{i}^{\mathbf{U}}$ does not incorporate the scaling term present in Eq. (6). Consequently, when different rows are quantized using the same bit-width, there is no significant variation in error. This uniformity arises from the fact that the Hessian matrices for different rows of $\mathbf{U}$ are identical. Thus, unlike $\mathbf{V}$, there is no necessity to employ mixed precision when quantizing different rows of $\mathbf{U}$.
Therefore, PrinMix determines the mixed-precision quantization schedule based on $\mathbf{V}$, and then applies the same schedule to $\mathbf{U}$ for simplicity. Specifically, PrinMix quantizes $\mathbf{U}$ using a column-wise mixed-precision schedule, where the $i^{\mathrm{th}}$ column of $\mathbf{U}$ adopts the same bit-width as the $i^{\mathrm{th}}$ row of $\mathbf{V}$ as they correspond to the same singular value. Notably, PrinMix exhibits insensitivity to column-wise precision schedules, since GPTQ compensates for quantization-induced errors in the column direction by adjusting the unquantized weights during the quantization process. This compensation, however, does not occur between different rows, as different rows are independently quantized in GPTQ. This further underscores the importance of discussing row-wise mixed precision strategies aimed at minimizing the quantization error of $\mathbf{V}$. In Appendix C.1, we further demonstrate experimentally that applying the same mixed-precision quantization strategy to both $\mathbf{V}$ and $\mathbf{U}$ yields satisfactory performance.
@@
In Eq. (6), we quantize $\mathbf{U}$ to reconstruct the target $\mathbf{U\Sigma}\hat{\mathbf{V}}X$, which deviates from the initial target $\mathbf{U\Sigma V}X$. This deviation can negatively impact the performance of the quantized model. A straightforward approach to address this issue is to directly replace the reconstruction target with $\mathbf{U\Sigma V}X$; however, this would inhibit the application of GPTQ for quantization, as it breaks the premise that GPTQ relies solely on second-order terms. Therefore, we propose a method termed “Reconstruction Target Correction” (RTC) to reduce the bias by transforming $\mathbf{U\Sigma\hat{V}}X$ in Eq. (6) to $\mathbf{\tilde{U}\Sigma\hat{V}}X$, where $\mathbf{\tilde{U}}$ is derived from the following equation:
@@
See Appendix A.3 for detailed derivations. In summary, prior to quantizing $\mathbf{U}$, we first update $\mathbf{U}$ to $\mathbf{\tilde{U}}$ using Eq. (7). Subsequently, we perform quantization by minimizing $\|\mathbf{\tilde{U}\Sigma\hat{V}}X-\mathbf{\hat{U}\Sigma\hat{V}}X\|_{F}^{2}$. This approach aims to ensure that the reconstruction target closely approximates the original, without compromising the application of GPTQ for quantization.
@@ 3.2 Optimization Problem Modeling
$\min_{\mathcal{S}}\sum_{i}\mathbb{E}^{\mathbf{V}}_{i}\mathcal{S}_{i}^{T}\ \ \ \ \text{(Total quantization error)}$ (8)
s.t. $\sum_{i}\mathcal{S}_{i}B\leq G_{b}(h_{\mathrm{in}}\cdot h_{\mathrm{out}})\ \ \ \ \text{(Bit budget constraint)}$
$\mathrm{sum}(S_{i})=1\ \ \ \ \text{(One-hot vector constraint)}$
$S_{i}-f\leq 0\ \ \ \ \text{(Bit-width selection constraint)}$
$\mathrm{sum}(f)\leq f_{\text{max}}\ \ \ \ \text{(Bit-width number constraint)}$
@@
[…]
By solving Eq. (8), we obtain an optimal mixed-precision scheme that minimizes the error while satisfying predefined bit budget constraints. This allows us to derive task-specific mixed-precision quantization strategies which balance the “scaling” and “difference” terms, leading to improved performance across various tasks.
Table 2: Comparison of PrinMix and baselines on various tasks across 7B-sized models. We report the results in the format “mean(std)” with three runs for Delta-CoMe and PrinMix.
@@
PrinMix 1/16 82.7(0.83) 36.7(3.35) 77.7(1.03) 94.6(0.51) 85.6(0.35) 83.1(0.25) 52.4(2.30) 79.4(0.83) 74.0
Table 3: Comparison of PrinMix and baselines on various tasks across 13-14B-sized models. We report the results in the format “mean(std)” with three runs for Delta-CoMe and PrinMix.
@@
PrinMix 1/16 80.2(2.09) 31.1(3.81) 21.7(0.64) 71.2(0.26) 91.5(0.60) 86.9(0.12) 62.7(0.04) 72.1(0.18) 64.7
4 Experiments
4.1 Experiment Setup
Evaluation Tasks
We evaluate our methods on four distinct tasks: reasoning, math, code generation, and multi-modal. These tasks encompass a vast array of current directions based on fine-tuning with open-source LLMs. Reasoning: We use the Math500 and AIME2024 datasets as the test set. Math: We use the GSM8K  (Cobbe et al., 2021) and Math500 (Lightman et al., 2023) datasets as the test set. Code Generation: We use HumanEval  (Chen et al., 2021) and MBPP  (Austin et al., 2021) as the test set. Multi-Modal: We utilize the GQA  (Hudson and Manning, 2019) and the image part of ScienceQA (Lu et al., 2022) datasets. Please refer to Appendix B.1 for more details.
Models
To ensure a comprehensive comparison, we evaluate both 7B and 13-14B models across the four tasks with various backbones, as shown in Table 1. During inference, we employ a greedy search strategy. Each model is compressed by a factor of 16 following (Ping et al., 2024) ($\alpha=1/16$). In Appendix C.2, we report the performance of PrinMix with different compression ratios.
Calibration Dataset
Following Delta-CoMe (Ping et al., 2024), PrinMix randomly samples 128 examples, each containing 2048 tokens, from the C4 training set as the calibration dataset. In Appendix C.3, we analyze the sensitivity of the calibration dataset by varying its domains and the number of examples. The results show that PrinMix performs consistently well across different calibration dataset configurations.
Baselines
We compare PrinMix with three baselines: SVD-based low-rank compression (Ryu et al., 2023), BitDelta  (Liu et al., 2024), and Delta-CoMe  (Ping et al., 2024). All methods are evaluated using NVIDIA L20 GPUs.
4.2 Main Results
Tables 2 and 3 present the results of PrinMix on both the 7B and 13-14B models across four tasks, in comparison to the baselines. Notably, PrinMix demonstrates superior overall performance on both the 7B and 13-14B models, surpassing the best baseline, Delta-CoMe, by an average of 2.9% and 2.2%, respectively.